OpenStudy (anonymous):

Given no other restrictions, what are the domain and range of the following function? f(x) = x^2 - 2x + 1

OpenStudy (hitaro9):

So, you know that the domain is going to be -infinity to infinity

OpenStudy (hitaro9):

Because you can plug any number in to x and get a result.

OpenStudy (agent0smith):

^so many students don't understand that concept, though.

OpenStudy (hitaro9):

You also know right off the bat that its an upward parabola, so it's going to go up, but it's going to have a minimum

OpenStudy (anonymous):

Thank you!!

OpenStudy (hitaro9):

The hard part is finding out what that minimum is.

OpenStudy (hitaro9):

But if you know how to graph a parabola, you know that the minimum is at 0

OpenStudy (hitaro9):

So the range is 0 to infinity.

OpenStudy (agent0smith):

^ the minimum of f(x) = x^2 - 2x + 1 will not be y = 0 First: Find the x-coordinate of the vertex, using x=-b/2a

OpenStudy (agent0smith):

The plug that into the function to find the minimum y value

OpenStudy (hitaro9):

Hmm

OpenStudy (hitaro9):

-(-2)/2(1)

OpenStudy (hitaro9):

Gives you 1

OpenStudy (hitaro9):

And at 1, (1)^2 -2(1) + 1

OpenStudy (hitaro9):

Gives you 0

OpenStudy (hitaro9):

So I'm pretty sure my range was right?

OpenStudy (agent0smith):

Turns out it was, but by chance (you didn't know above that it was).

OpenStudy (agent0smith):

"But if you know how to graph a parabola, you know that the minimum is at 0" this statement... just because it's a parabola doesn't mean the min is 0

OpenStudy (hitaro9):

Eh. I just assumed that they knew how to find the vertex of a parabola.